OpenStudy (anonymous):

@Mathematics Three vertices of a parallelogram are (2,7), (6,1) and (9,6) . Find the fourth vertex.

OpenStudy (anonymous):

the vector from (6,1) to (9,6) and the vector from (2,7) to (x,y) are the same vector first vector is (-6,-1) + (9,6) = (3,5) hence second vector, which in terms of x and y is (-2,-7) + (x,y) = (x - 2, y - 7) is equal to (3,5) hence x - 2 = 3, x = 5 hence y - 7 = 5, y = 12 hence other vertex is (5, 12)

OpenStudy (anonymous):

why did (6,1) became (-6,-1)? :D

OpenStudy (anonymous):

have you done vectors before?

OpenStudy (phi):



OpenStudy (phi):

a vector is head-tail, so james found (9,6) - (6,1) for the vector going from point 6,1 to point 9,6 He then added this vector to point 2,7 to get the 4th point

OpenStudy (anonymous):

Oh! Now I remember a bit. Thanks james and phi :)